OpenStudy (anonymous):

HOW TO SOLVE THESE TYPE OF PROBLEM: 5^x + 5x =1

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

there is no analytical approach !!

OpenStudy (lgbasallote):

there's only one possible answer =_=

OpenStudy (anonymous):

0?

OpenStudy (lgbasallote):

yes lol

OpenStudy (lgbasallote):

wonder how to solve ti though

OpenStudy (anonymous):

Actually I need to solve this: 400 + 400n + 10001 ( 1.0001^n - 1) > 29000

OpenStudy (lgbasallote):

same shiz

OpenStudy (anonymous):

LHS is an increasing function so just trial and error...

OpenStudy (anonymous):

because there is no analytical way to solve such equations

OpenStudy (anonymous):

http://openstudy.com/users/shfreeman#/updates/4f56c07ae4b0073f91c7fa3f scroll to the bottom of this question

OpenStudy (anonymous):

yeah just numerical methods..

OpenStudy (anonymous):

for equations like this\[x=\sin x\]\[3^x+2x=9\]\[x-e^x=\sin x\]\[\ln x=x+1\]...

OpenStudy (anonymous):

400 + 400n + 10001 ( 1.0001^n - 1) > 29000 BUT THIS CAN BE SOLVED

OpenStudy (anonymous):

WITHOUT TRIAL AND ERROR METHOD

OpenStudy (anonymous):

exact value of n?

OpenStudy (anonymous):

YEP

OpenStudy (anonymous):

i mean for real numbers n

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Oh..... I guess I did a mistake

OpenStudy (anonymous):

Sorry guys...... And thanx @mukushla and all who gave their precious time

OpenStudy (anonymous):

anytime saura